The difference between two numbers is 8, if 2 be added to the greater the result will be 3 times the smaller. find the numbers
Answer by macston(5194)

.
x=larger number; y=smaller number
.
x-y=8
x=y+8 Use this to substitute for x.
.
x+2=3y Substitute for x from above.
(y+8)+2=3y
y+10=3y Subtract y from each side.
10=2y Divide each side by 2.
5=y ANSWER 1: the smaller number is 5.
.
x=y+8=5+8=13 ANSWER 2: The larger number is 13.
.
CHECK:
x+2=3y
13+2=3(5)
15=15